# Towards Affordance-Aware Robotic Dexterous Grasping with Human-like Priors

### v3

@@ Abstract
Figure 1: Performance comparision among UniDexGrasp (Xu et al. 2023), UniDexGrasp++ (Wan et al. 2023), and our AffordDex, on the vision-based setting. we report human-likeness score (HLS) and affordance score (AS) across seen objects, unseen objects, and unseen categories. We also present a qualitative comparison, where AffordDex performs natural and safe grasping by avoiding the blade.
@@ 1 Introduction
Dexterous grasping, as a foundational capability for robotic manipulation, has garnered significant attention from both academia and industry (Zhao et al. 2024b). Compared to simpler end-effectors (e.g., parallel jaws, vacuum grippers), five-fingered dexterous hands closely resemble human hand structure, providing substantially enhanced flexibility, precision, and task adaptability (Zhong et al. 2025). Furthermore, anthropomorphic robots expedite the collection of rich human demonstration data via teleoperation (Li et al. 2025a). Consequently, this synergy has fueled rapid progress, with recent algorithms achieving high success rates in generalizing grasps to novel objects (Fang et al. 2022, 2020; Gou et al. 2021; Wang et al. 2021; Xu et al. 2023; Wan et al. 2023).
Due to the high degrees of freedom (DOFs) of dexterous hands, traditional motion planning-based methods (Andrews and Kry 2013; Bai and Liu 2014) struggle to handle such complex hand joint movements. Recent advancements in reinforcement learning (RL) (Wan et al. 2023; Mandikal and Grauman 2022; Christen et al. 2022; Nagabandi et al. 2020; Mandikal and Grauman 2021) have shown promising results in complex dexterous manipulation. However, the goal of grasping is not merely to lift an object. It involves alignment with human intent and preparation for subsequent manipulation tasks, such as avoiding the blade of a knife or preparing to open a bottle cap. Existing methods, while focused on low-level grasp stability metrics, largely overlook this crucial synthesis of affordance-aware positioning and human-like kinematics, limiting their utility in real-world, multi-step manipulation scenarios.
@@ 2.1 Dexterous Grasping
Robotic grasping (Fang et al. 2022, 2020; Gou et al. 2021; Wang et al. 2021) has been a longstanding research, aiming to enable robots to interact with objects reliably and adaptively. While significant advances have been made with simple parallel-jaw grippers (Fang et al. 2020; Mahler et al. 2019), their limited dexterity restricts adaptability to objects with intricate geometries. Dexterous, multi-fingered hands (Xu et al. 2023; Wan et al. 2023) offer a solution but pose a severe control challenge for traditional analytical methods (Bai and Liu 2014; Liu et al. 2021), motivating the shift towards learning-based approaches.
One paradigm decouples the grasping process into static grasp pose generation followed by a dynamic grasping through trajectory planning or goal-conditioned reinforcement learning (RL) (Wan et al. 2023; Christen et al. 2022; Wang et al. 2025). For example, UniDexGrasp++ proposes geometry-aware curriculum learning and leverages the geometry feature for RL. However, these RL-based methods may produce physically unrealistic joint configurations. An alternative paradigm directly learns the entire grasping trajectory through expert demonstrations from humans or reinforcement learning agents (Xu et al. 2023; Liu et al. 2024; Huang et al. 2023; Lu et al. 2024; Zhang et al. 2024b, a). These approaches tend to achieve more natural motions but suffer from poor generalization to novel objects due to the limited diversity of demonstrations and inherent policy constraints. To address these failures, our AffordDex combines strong, human-derived motion priors to ensure natural movement with affordance-based guidance to achieve robust generalization, resulting in a policy that is both natural and functionally effective across a wide range of objects.
@@ 2.2 Affordance Prediction
Affordance defines the action possibilities an object offers to an agent (Gibson 2014). In robotics, this translates to identifying object regions suitable for specific interactions, such as grasping, pushing, or lifting. Predicting such affordances is therefore critical for advanced visual understanding and robotic manipulation, as evidenced by extensive research (Li et al. 2019; Cao et al. 2020; Corona et al. 2020; Jiang et al. 2021; Lu et al. 2025; Shao et al. 2025). While pioneering works like GanHand (Corona et al. 2020) introduced generative models for multi-object on-table grasps, and GEAL (Lu et al. 2025) pioneered a dual-branch architecture for cross-modal (3D point cloud to 2D) representation learning, their learned affordances are often task- or category-specific. This inherent specialization limits their ability to generalize to novel objects or adapt to different downstream manipulation requirements. By contrast, humans exhibit exceptional proficiency in inferring universal affordances from visual cues (Zhao et al. 2024a). Inspired by this capability, AffordDex learns to infer functional affordances directly from multi-view rendered images of 3D objects, enabling a generalizable grasping policy that is not constrained to specific object categories or predefined tasks.
@@ 3 Methodology
To generate grasps with affordance-aware positioning and human-like kinematics, crucial for facilitating downstream manipulation, we propose a novel two-stage framework. The first stage establishes a strong human motion prior by pre-training a base policy $\pi^{H}$, on a large-scale human motion dataset (Zhan et al. 2024) via imitation learning. This constrains the policy to a manifold of natural, human-like movements. In the second stage, we freeze the weights of $\pi^{H}$ and train a lightweight residual module via reinforcement learning (RL) to adapt these general motions to specific object interactions. This RL refinement stage is critically guided by two components: our Negative Affordance-aware Segmentation (NAA) module, which provides explicit constraints on where not to touch an object, and a teacher-student distillation framework that leverages privileged state information to significantly boost the final policy’s performance. An overview of our method is illustrated in Fig. 2.
@@ 3.1 Human Hand Trajectory Imitating
The finger imitation reward $r^{H}_{\text{finger}}$ encourages the dexterous hand to closely track the reference finger poses from human hand dataset. Following (Li et al. 2025b), we define this reward based on the distance between the corresponding keypoints $F$ on the robot dexterous hand and the MANO hand. The reward at time $t$ is formulated as:
@@ 3.2 Negative Affordance-aware Segmentation
A significant limitation of prior work in grasp synthesis (Xu et al. 2023; Wan et al. 2023; Zhong et al. 2025), is its neglect of the semantic and functional context of the interaction. A classic example is a knife: while its blade is geometrically stable for grasping, any such grasp is functionally incorrect and unsafe. To address limitation, we introduce the Negative Affordance-aware Segmentation (NAA) module to incorporate negative affordances—reasoning about which parts of an object should not be touched. The proposed NAA has the ability to operate in an open-vocabulary manner by harnessing the rich world knowledge embedded in Vision-Language Models (VLMs) (Radford et al. 2021; Achiam et al. 2023), automatically benefiting from future progress in foundation models. This ensures that the generated grasps are not only geometrically stable but also semantically coherent and task-aware.
VLMs struggle to interpret non-textured 3D meshes, as these models primarily rely on rich visual cues learned from images. To bridge this gap, we first apply procedural texturing to the raw meshes by (Zhang et al. 2024c), which generates semantically plausible textures based on geometric analysis, ensuring robustness across different object shapes. Next, we render the textured object from six cardinal directions to create a multi-view image set $I$ as a holistic visual representation. While this may not capture all concavities in highly complex objects, we found it provides a sufficient basis for affordance prediction for objects in the benchmark datasets, representing a practical trade-off between coverage and computational cost. We then query GPT-4V (Achiam et al. 2023) to elicit a detailed description of the object’s negative affordances.
VLMs (Radford et al. 2021) and Multimodal Large Language Models (MLLMs) (Achiam et al. 2023) excel at image-level understanding but struggle with the fine-grained spatial localization required for segmentation. To solve this, instead of asking CLIP (Radford et al. 2021) to find “blade part” from the image, we turn the segmentation task into a much simpler classification task. We generate a set of precise object-part masks $M_{i}$ and use them as a visual prompt to let CLIP identify which mask in $M_{i}$ has the highest semantic similarity to the textual description “blade part”.
Specifically, for each image $I_{i}\in I$, we prompt Segment Anything Model (SAM) (Kirillov et al. 2023) with a dense grid of points $G$ overlaid on $I_{i}$, which prompts SAM to perform an exhaustive segmentation, identifying all potential objects and parts. The resulting collection of masks is then refined using Non-Maximum Suppression (NMS) to eliminate duplicates, yielding a clean candidate masks set $M_{i}$:
@@
For each mask $M_{i}^{j}\in M_{i}$, we generate a visually prompted image $I_{i}^{j}$ by blurring regions outside the mask with Gaussian filter following (Yang et al. 2023). The prompted image set $\{I_{i}^{j}\}$ is then passed to CLIP along with the text query to compute a similarity score for each image-text pair. The mask with the highest similarity score is selected as the final segmentation mask. The mask is then projected into 3D space to segment the corresponding regions of the object’s point cloud to get the negative affordance $N_{t}$, as shown in Fig. 3.
@@
Figure 4: Qualitative Comparison on UniDexGrasp (Xu et al. 2023) and OakInk2 (Zhan et al. 2024). A comparison of grasps generated by our AffordDex with several baselines, including UniDexGrasp (Xu et al. 2023), UniDexGrasp++ (Wan et al. 2023), and DexGrasp Anything (Zhong et al. 2025).
@@ 3.3 Affordance-aware Residual Learning
The negative affordance predicted from proposed NAA, we use a residual module $\mathcal{R}$ to refine the pre-trained policy $\pi^{H}$. Since visual pose estimation is inherently less precise than using privileged state information, directly training an effective vision-based policy can be challenging. Therefore, we first train a state-based teacher policy $\pi^{T}$ which can access the ground-truth states of the environment, such as object states, to learn residual actions to refine the initial actions predicted by $\pi^{H}$. Once the teacher policy $\pi^{T}$ finishes training, we use an imitation learning algorithm, DAgger (Ross, Gordon, and Bagnell 2011), to distill $\pi^{T}$ to a vision-based student policy $\pi^{S}$ that can access oracle information and let policy help and ease the vision-based policy learning.
State-based teacher policy. In this stage, the inputs are robot state $R_{t}$, object state $O_{t}$, the scene point clouds $P_{t}$, and predicted negative affordance $N_{t}$. Here the scene point cloud is fused by multi-view depth cameras. Our goal is to learn residual actions $\Delta a_{t}=\pi^{T}(S^{T}_{t})$ with predicted negative affordance by PPO (Schulman et al. 2017). The final action is computed with an element-wise addition:
@@
Vision-based student policy. For vision-based policy, we only allow it to access information available in the real world, including robot state $R_{t}$, the scene points clouds $P_{t}$, and predicted negative affordance $N_{t}$. Then, we distill the teacher policy $\pi^{T}$ into vision-based student policy $\pi^{S}$ using DAgger (Ross, Gordon, and Bagnell 2011), i.e.,
@@ 4.1 Datasets
UniDexGrasp (Xu et al. 2023). This dataset contains 3165 different object instances spanning 133 categories. Evaluation is conducted on these 3,200 seen objects, as well as on 140 unseen objects from seen categories and 100 unseen objects from unseen categories. Each environment is randomly initialized with one object and its initial pose, and the environment consists of a panoramic 3D point cloud $P_{t}$ captured from the fixed cameras for vision-based policy learning.
OakInk2 (Zhan et al. 2024). This dataset record the manipularion processes with pose and shape of the human upper-body and objects. We pre-train our $\pi^{H}$ using about 2,200 right hand manipulation sequences in this dataset. Also we employ objects in OakInk2 to evaluate the generalization capabilities for grasping.
Table 1: Quantitative comparisons on UniDexGrasp (Xu et al. 2023) and OakInk2 (Zhan et al. 2024). HLSHLS denotes Human-likeness Score, while ASAS is Affordance Score.
@@
PPO (Schulman et al. 2017) 24.3 - - 20.9 - - 17.2 - - - - -
DAPG (Rajeswaran et al. 2017) 20.8 - - 15.3 - - 11.1 - - - - -
GSL (Jia et al. 2022) 57.3 - - 54.1 - - 50.9 - - - - -
ILAD (Wu, Wang, and Wang 2023) 31.9 - - 26.4 - - 23.1 - - - - -
UniDexGrasp (Xu et al. 2023) 79.4 6.9 12 74.3 6.4 15 70.8 6.3 18 68.4 5.9 18
UniDexGrasp++ (Wan et al. 2023) 87.9 5.4 28 84.3 5.2 26 83.1 5.0 27 79.6 4.9 28
DexGrasp Anything (Zhong et al. 2025) 71.2 - 20 69.1 - 18 67.3 - 22 65.9 - 24
@@
PPO (Schulman et al. 2017) 20.6 - - 17.2 - - 15.0 - - - - -
DAPG (Rajeswaran et al. 2017) 20.8 - - 15.3 - - 11.1 - - - - -
GSL (Jia et al. 2022) 54.1 - - 50.2 - - 44.8 - - - - -
ILAD (Wu, Wang, and Wang 2023) 27.6 - - 23.2 - - 20.0 - - - - -
UniDexGrasp (Xu et al. 2023) 73.7 6.2 16 68.6 6.1 18 65.1 6.0 17 62.8 5.6 20
UniDexGrasp++ (Wan et al. 2023) 85.4 5.4 29 79.6 5.1 25 76.7 4.8 28 74.4 4.7 29
@@
Table 2: Ablation Study on UniDexGrasp (Xu et al. 2023) in Seen Object.
@@
Table 3: Results on UniDexGrasp (Xu et al. 2023) in Seen Object in state-based setting.

### v4

@@ Abstract
Figure 1: Performance comparision among UniDexGrasp (xu2023unidexgrasp), UniDexGrasp++ (wan2023unidexgrasp++), and our AffordDex, on the vision-based setting. we report human-likeness score (HLS) and affordance score (AS) across seen objects, unseen objects, and unseen categories. We also present a qualitative comparison, where AffordDex performs natural and safe grasping by avoiding the blade.
@@ 1 Introduction
Dexterous grasping, as a foundational capability for robotic manipulation, has garnered significant attention from both academia and industry (zhao2024sg; zhao2025high). Compared to simpler end-effectors (e.g., parallel jaws, vacuum grippers), five-fingered dexterous hands closely resemble human hand structure, providing substantially enhanced flexibility, precision, and task adaptability (zhong2025dexgrasp). Furthermore, anthropomorphic robots expedite the collection of rich human demonstration data via teleoperation (li2025teleopbench). Consequently, this synergy has fueled rapid progress, with recent algorithms achieving high success rates in generalizing grasps to novel objects (fang2022transcg; fang2020graspnet; gou2021rgb; wang2021graspness; xu2023unidexgrasp; wan2023unidexgrasp++).
Due to the high degrees of freedom (DOFs) of dexterous hands, traditional motion planning-based methods (andrews2013goal; bai2014dexterous) struggle to handle such complex hand joint movements. Recent advancements in reinforcement learning (RL) (wan2023unidexgrasp++; mandikal2022dexvip; christen2022d; nagabandi2020deep; mandikal2021learning) have shown promising results in complex dexterous manipulation. However, the goal of grasping is not merely to lift an object. It involves alignment with human intent and preparation for subsequent manipulation tasks, such as avoiding the blade of a knife or preparing to open a bottle cap. Existing methods, while focused on low-level grasp stability metrics, largely overlook this crucial synthesis of affordance-aware positioning and human-like kinematics, limiting their utility in real-world, multi-step manipulation scenarios.
@@ 2.1 Dexterous Grasping
Robotic grasping (fang2022transcg; fang2020graspnet; gou2021rgb; wang2021graspness) has been a longstanding research, aiming to enable robots to interact with objects reliably and adaptively. While significant advances have been made with simple parallel-jaw grippers (fang2020graspnet; mahler2019learning), their limited dexterity restricts adaptability to objects with intricate geometries. Dexterous, multi-fingered hands (xu2023unidexgrasp; wan2023unidexgrasp++) offer a solution but pose a severe control challenge for traditional analytical methods (bai2014dexterous; liu2021synthesizing), motivating the shift towards learning-based approaches.
One paradigm decouples the grasping process into static grasp pose generation followed by a dynamic grasping through trajectory planning or goal-conditioned reinforcement learning (RL) (wan2023unidexgrasp++; christen2022d; wang2025unigrasptransformer). For example, UniDexGrasp++ proposes geometry-aware curriculum learning and leverages the geometry feature for RL. However, these RL-based methods may produce physically unrealistic joint configurations. An alternative paradigm directly learns the entire grasping trajectory through expert demonstrations from humans or reinforcement learning agents (xu2023unidexgrasp; liu2024realdex; huang2023diffusion; lu2024ugg; zhang2024artigrasp; zhang2024graspxl). These approaches tend to achieve more natural motions but suffer from poor generalization to novel objects due to the limited diversity of demonstrations and inherent policy constraints. To address these failures, our AffordDex combines strong, human-derived motion priors to ensure natural movement with affordance-based guidance to achieve robust generalization, resulting in a policy that is both natural and functionally effective across a wide range of objects.
@@ 2.2 Affordance Prediction
Affordance defines the action possibilities an object offers to an agent (gibson2014theory). In robotics, this translates to identifying object regions suitable for specific interactions, such as grasping, pushing, or lifting. Predicting such affordances is therefore critical for advanced visual understanding and robotic manipulation, as evidenced by extensive research (li2019putting; cao2020long; corona2020ganhand; jiang2021hand; lu2025geal; shao2025great). While pioneering works like GanHand (corona2020ganhand) introduced generative models for multi-object on-table grasps, and GEAL (lu2025geal) pioneered a dual-branch architecture for cross-modal (3D point cloud to 2D) representation learning, their learned affordances are often task- or category-specific. This inherent specialization limits their ability to generalize to novel objects or adapt to different downstream manipulation requirements. By contrast, humans exhibit exceptional proficiency in inferring universal affordances from visual cues (zhao2024automated). Inspired by this capability, AffordDex learns to infer functional affordances directly from multi-view rendered images of 3D objects, enabling a generalizable grasping policy that is not constrained to specific object categories or predefined tasks.
@@ 3 Methodology
To generate grasps with affordance-aware positioning and human-like kinematics, crucial for facilitating downstream manipulation, we propose a novel two-stage framework. The first stage establishes a strong human motion prior by pre-training a base policy $\pi^{H}$, on a large-scale human motion dataset (zhan2024oakink2) via imitation learning. This constrains the policy to a manifold of natural, human-like movements. In the second stage, we freeze the weights of $\pi^{H}$ and train a lightweight residual module via reinforcement learning (RL) to adapt these general motions to specific object interactions. This RL refinement stage is critically guided by two components: our Negative Affordance-aware Segmentation (NAA) module, which provides explicit constraints on where not to touch an object, and a teacher-student distillation framework that leverages privileged state information to significantly boost the final policy’s performance. An overview of our method is illustrated in Fig. 2.
@@ 3.1 Human Hand Trajectory Imitating
The finger imitation reward $r^{H}_{\text{finger}}$ encourages the dexterous hand to closely track the reference finger poses from human hand dataset. Following (li2025maniptrans), we define this reward based on the distance between the corresponding keypoints $F$ on the robot dexterous hand and the MANO hand. The reward at time $t$ is formulated as:
@@ 3.2 Negative Affordance-aware Segmentation
A significant limitation of prior work in grasp synthesis (xu2023unidexgrasp; wan2023unidexgrasp++; zhong2025dexgrasp), is its neglect of the semantic and functional context of the interaction. A classic example is a knife: while its blade is geometrically stable for grasping, any such grasp is functionally incorrect and unsafe. To address limitation, we introduce the Negative Affordance-aware Segmentation (NAA) module to incorporate negative affordances—reasoning about which parts of an object should not be touched. The proposed NAA has the ability to operate in an open-vocabulary manner by harnessing the rich world knowledge embedded in Vision-Language Models (VLMs) (radford2021learning; achiam2023gpt), automatically benefiting from future progress in foundation models. This ensures that the generated grasps are not only geometrically stable but also semantically coherent and task-aware.
VLMs struggle to interpret non-textured 3D meshes, as these models primarily rely on rich visual cues learned from images. To bridge this gap, we first apply procedural texturing to the raw meshes by (zhang2024texpainter), which generates semantically plausible textures based on geometric analysis, ensuring robustness across different object shapes. Next, we render the textured object from six cardinal directions to create a multi-view image set $I$ as a holistic visual representation. While this may not capture all concavities in highly complex objects, we found it provides a sufficient basis for affordance prediction for objects in the benchmark datasets, representing a practical trade-off between coverage and computational cost. We then query GPT-4V (achiam2023gpt) to elicit a detailed description of the object’s negative affordances.
VLMs (radford2021learning) and Multimodal Large Language Models (MLLMs) (achiam2023gpt) excel at image-level understanding but struggle with the fine-grained spatial localization required for segmentation. To solve this, instead of asking CLIP (radford2021learning) to find “blade part” from the image, we turn the segmentation task into a much simpler classification task. We generate a set of precise object-part masks $M_{i}$ and use them as a visual prompt to let CLIP identify which mask in $M_{i}$ has the highest semantic similarity to the textual description “blade part”.
Specifically, for each image $I_{i}\in I$, we prompt Segment Anything Model (SAM) (kirillov2023segment) with a dense grid of points $G$ overlaid on $I_{i}$, which prompts SAM to perform an exhaustive segmentation, identifying all potential objects and parts. The resulting collection of masks is then refined using Non-Maximum Suppression (NMS) to eliminate duplicates, yielding a clean candidate masks set $M_{i}$:
@@
For each mask $M_{i}^{j}\in M_{i}$, we generate a visually prompted image $I_{i}^{j}$ by blurring regions outside the mask with Gaussian filter following (yang2023fine). The prompted image set $\{I_{i}^{j}\}$ is then passed to CLIP along with the text query to compute a similarity score for each image-text pair. The mask with the highest similarity score is selected as the final segmentation mask. The mask is then projected into 3D space to segment the corresponding regions of the object’s point cloud to get the negative affordance $N_{t}$, as shown in Fig. 3. Our NAA is an offline, one-time process. About 160 seconds per object on an RTX 4090 is a reasonable one-time trade-off.
@@
Figure 4: Qualitative Comparison on UniDexGrasp (xu2023unidexgrasp) and OakInk2 (zhan2024oakink2). A comparison of grasps generated by our AffordDex with several baselines, including UniDexGrasp (xu2023unidexgrasp), UniDexGrasp++ (wan2023unidexgrasp++), and DexGrasp Anything (zhong2025dexgrasp).
@@ 3.3 Affordance-aware Residual Learning
The negative affordance predicted from proposed NAA, we use a residual module $\mathcal{R}$ to refine the pre-trained policy $\pi^{H}$. Since visual pose estimation is inherently less precise than using privileged state information, directly training an effective vision-based policy can be challenging. Therefore, we first train a state-based teacher policy $\pi^{T}$ which can access the ground-truth states of the environment, such as object states, to learn residual actions to refine the initial actions predicted by $\pi^{H}$. Once the teacher policy $\pi^{T}$ finishes training, we use an imitation learning algorithm, DAgger (ross2011reduction), to distill $\pi^{T}$ to a vision-based student policy $\pi^{S}$ that can access oracle information and let policy help and ease the vision-based policy learning.
State-based teacher policy. In this stage, the inputs are robot state $R_{t}$, object state $O_{t}$, the scene point clouds $P_{t}$, and predicted negative affordance $N_{t}$. Here the scene point cloud is fused by multi-view depth cameras. Our goal is to learn residual actions $\Delta a_{t}=\pi^{T}(S^{T}_{t})$ with predicted negative affordance by PPO (schulman2017proximal). The final action is computed with an element-wise addition:
@@
Vision-based student policy. For vision-based policy, we only allow it to access information available in the real world, including robot state $R_{t}$, the scene points clouds $P_{t}$, and predicted negative affordance $N_{t}$. Then, we distill the teacher policy $\pi^{T}$ into vision-based student policy $\pi^{S}$ using DAgger (ross2011reduction), i.e.,
@@ 4.1 Datasets
UniDexGrasp (xu2023unidexgrasp). This dataset contains 3165 different object instances spanning 133 categories. Evaluation is conducted on these 3,200 seen objects, as well as on 140 unseen objects from seen categories and 100 unseen objects from unseen categories. Each environment is randomly initialized with one object and its initial pose, and the environment consists of a panoramic 3D point cloud $P_{t}$ captured from the fixed cameras for vision-based policy learning.
OakInk2 (zhan2024oakink2). This dataset record the manipularion processes with pose and shape of the human upper-body and objects. We pre-train our $\pi^{H}$ using about 2,200 right hand manipulation sequences in this dataset. Also we employ objects in OakInk2 to evaluate the generalization capabilities for grasping.
Table 1: Quantitative comparisons on UniDexGrasp (xu2023unidexgrasp) and OakInk2 (zhan2024oakink2). HLSHLS denotes Human-likeness Score, while ASAS is Affordance Score.
@@
PPO (schulman2017proximal) 24.3 - - 20.9 - - 17.2 - - - - -
DAPG (rajeswaran2017learning) 20.8 - - 15.3 - - 11.1 - - - - -
GSL (jia2022improving) 57.3 - - 54.1 - - 50.9 - - - - -
ILAD (wu2023learning) 31.9 - - 26.4 - - 23.1 - - - - -
UniDexGrasp (xu2023unidexgrasp) 79.4 6.9 12 74.3 6.4 15 70.8 6.3 18 68.4 5.9 18
UniDexGrasp++ (wan2023unidexgrasp++) 87.9 5.4 28 84.3 5.2 26 83.1 5.0 27 79.6 4.9 28
DexGrasp Anything (zhong2025dexgrasp) 71.2 - 20 69.1 - 18 67.3 - 22 65.9 - 24
@@
PPO (schulman2017proximal) 20.6 - - 17.2 - - 15.0 - - - - -
DAPG (rajeswaran2017learning) 20.8 - - 15.3 - - 11.1 - - - - -
GSL (jia2022improving) 54.1 - - 50.2 - - 44.8 - - - - -
ILAD (wu2023learning) 27.6 - - 23.2 - - 20.0 - - - - -
UniDexGrasp (xu2023unidexgrasp) 73.7 6.2 16 68.6 6.1 18 65.1 6.0 17 62.8 5.6 20
UniDexGrasp++ (wan2023unidexgrasp++) 85.4 5.4 29 79.6 5.1 25 76.7 4.8 28 74.4 4.7 29
@@
Table 2: Ablation Study on UniDexGrasp (xu2023unidexgrasp) in Seen Object.
@@
Table 3: Results on UniDexGrasp (xu2023unidexgrasp) in Seen Object in state-based setting.